# Rethinking Data Input for Point Cloud Upsampling

### v2

@@
Abstract.
Point cloud upsampling is crucial for tasks like 3D reconstruction. While existing methods rely on patch-based inputs, and there is no research discussing the differences and principles between point cloud model full input and patch based input. Ergo, we propose a novel approach using whole model inputs i.e. Average Segment input. Our experiments on PU1K and ABC datasets reveal that patch-based inputs consistently outperform whole model inputs. To understand this, we will delve into factors in feature extraction, and network architecture that influence upsampling results.
@@ Abstract.
ccs: Computing methodologies Computer vision
1. Introduction
@@
The improvement of 3D scanners in 3D sensors, as well as the rapid development of deep learning in computer vision, have led to a gradual shift in the popularity of computer vision technology from 2D image to 3D model. Especially in the direction of autonomous driving and robots that exchange with the physical world, there is a particular focus on three-dimensional vision (Endres et al., 2013; Roriz et al., 2021; Naseer et al., 2018; Zhou and Tuzel, 2018).
@@
Figure 1. The original Mesh models have been transformed into a point cloud through patch based sampling and the Average Segment method proposed in this article.
@@
Vis-a-vis other representations such as Mesh, point clouds are an effective way to represent three-dimensional shapes that are widely used in the aforementioned fields. In addition, the task of reverse engineering CAD models from 3D point clouds is receiving increasing attention (Alai, 2013). In the process of redesigning the model obtained from reverse engineering, although there are Application Programming Interface (API) available in CAD software (Kao and Lin, 1998), if these designs involve reverse engineering and highly customized and personalized designs, CAD APIs are powerless (Shatnawi et al., 2017). Therefore, the generation of model point clouds is crucial.
@@
The point cloud upsampling technology based on deep learning has developed with the development of deep learning, extracting multi-resolution features from the initial PU-Net (Yu et al., 2018) on the basis of pointnet++ (Qi et al., 2017). Then, multiple independent multi-layer perceptrons (MLPs) are introduced to generate the required dense point cloud. Moving on to MPU/3PU (Yifan et al., 2019), KNN is used for interpolation and EdgeConv (Wang et al., 2019) is used to extract features for hierarchical reconstruction through recursive upsampling. Implementing upsampling PU-GAN (Li et al., 2019) based on GAN (Goodfellow et al., 2020) generation architecture, guiding the generator to upsample and approximate the real point cloud. PU-GCN (Qian et al., 2021) modifies the upsampling module from simple duplicate to Nodeshuffle based on feature extraction in MPU, achieving very good results. And Transformer’s point cloud upsampling network (Qiu et al., 2022; Vaswani et al., 2017). There are also methods to create coarse point clouds and then refine them by combining global and local methods (Li et al., 2021).
@@
The above methods mainly utilize Euclidean spatial features at potential spatial points, upsampling from low resolution to high resolution. The inherent limitations have also been addressed by the work of upsampling point clouds at any scale (Zhao et al., 2022; He et al., 2023) and capturing local information (Kim et al., 2024).
@@
Nevertheless, these works still have limitations, which are on the training dataset. Existing point cloud upsampling work (Qian et al., 2021; Qiu et al., 2022; Zhao et al., 2022; He et al., 2023; Kim et al., 2024; Mao et al., 2022) without exception used the PU1K dataset (Qian et al., 2021), which was used based on the methods of MPU and PU-GAN (Yifan et al., 2019; Li et al., 2019). To obtain training data by randomly sampling the entire 3D model using Patch under conditions of limited computing power, especially limited graphics memory. Undoubtedly, Patch based data input can reduce memory consumption and computational complexity, but it can bring about boundary effects and continuity loss in reverse engineering. It is necessary to explore a new point cloud input method to avoid the above issues. Moreover, the existing point cloud upsampling work focuses on 3D models similar to the PU1K dataset, while neglecting the exploration of point cloud upsampling for parts and CAD. In addition to conducting experiments on the PU1K dataset, this article will also discuss the upsampling effect of point cloud models on workpieces on the ABC dataset (Koch et al., 2019).
@@
Based on the above issues, in order to explore why numerous point cloud sampling data are still based on Patch input, this paper proposes a new data input method. Unlike dividing point clouds according to Patch, this article takes the point cloud model uniformly and randomly divided and corresponds them one by one as input, which is named as Average Segment (AS). After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN (Qian et al., 2021) network, and cross validation was performed on both the PU1k and ABC datasets. And during the experiment, the encoder and decoder of PU-GCN (Qian et al., 2021) were ablated.
@@
To solve the above problems, this paper proposes a new data input method. Unlike dividing point clouds according to patches, this paper uniformly and randomly divides point cloud models and uses them as inputs one by one. After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN (Qian et al., 2021), and cross validation was performed on both the PU1k (Qian et al., 2021) and ABC datasets (Koch et al., 2019). And during the experiment, the encoder and decoder of PU-GCN (Qian et al., 2021) were ablated.
@@ 2.1. Overview
2.2. Data Input Based on Patch
@@
2.3. Data Input Based on Average Segment
@@ 3.1. Settings
  • PU1K: This is a point cloud upsampling dataset introduced in PU-GCN (Qian et al., 2021). Overall, PU1K consists of 1147 3D models, including 120 3D models from the PU-GAN dataset, as well as 900 different models collected from ShapeNetCore (Chang et al., 2015). The test set includes 27 models from PU-GAN and over 100 models from ShapeNetCore. Covering 50 object categories.
  • ABC Dataset: The ABC Dataset (Koch et al., 2019) is a dataset containing one million computer-aided design (CAD) models, aimed at promoting research on geometric deep learning methods and applications. It is currently one of the largest publicly available datasets for geometric deep learning research.
@@
Figure 2. The PU-GCN network structure used in the experiment is shown in the figure, where (a) is the original PU-GCN structure. (b), (c) and (d) are all architectures used in the ablation experiment to explore the modules that have the greatest impact on upsampling.
3.2. Evaluation Metrics
During the testing process, this article adopted the common practices used in previous work on point cloud upsampling (Qian et al., 2021; Qiu et al., 2022; Zhao et al., 2022; He et al., 2023; Kim et al., 2024; Mao et al., 2022). That is to say, this article will cut the input point cloud into multiple subset Patch and AS of all N points.
In the 4 times upsampling experiment of this article, three types of low resolution to high-resolution sampling were performed. Each test sample has a low resolution point cloud consisting of 256, 512 and 1048 points, and a high-resolution point cloud consisting of 1024,2048 and 8196 points. This article quantitatively evaluates the upsampling performance of PU-GCN under two methods based on three widely used indicators: (i) Chamfer distance (CD) (Hajdu et al., 2012), (ii) Hausdorff distance (HD) (Berger et al., 2013), and (iii) model parameter quantity.
3.3. Results
@@
Table 1. The data input based on the Patch method and AS method is obtained using PU-GCN in PU1K.
@@
Table 2. The data input based on the Patch method and AS method is obtained using PU-GCN in ABC10K.
@@ 3.4. Ablations
Table 3. The results of swapping test sets based on Patch method and AS method on two datasets.
@@
Table 4. Perform ablation experiments on different architectures of PU-GCN using two data input methods.
@@ 3.5. Visualizations
Figure 3. Visualization of PU-GCN with Different Architectures on the PU1K Dataset.
Figure 4. Visualization example on the ABC dataset.
4. Discussion
@@
5. Future Works
PU-GCN model used in this article as an example, the existing Patch based upsampling method is better than AS method we proposed. This may be due to the difference in dataset size between the two methods. In future work, we will delve deeper into the differences in interpretability between Patch and AS methods through representation learning methods. At the same time, due to the similarity between the use of point clouds and the completion of point clouds. We will refer to the method in article  (Xia et al., 2021) and use encoders for the data input of Patch and AS respectively, in order to merge global and edge features and achieve better results.
6. Conclusion

### v3

@@
Abstract
Point cloud upsampling is crucial for tasks like 3D reconstruction. While existing methods rely on patch-based inputs, and there is no research discussing the differences and principles between point cloud model full input and patch based input. Ergo, we propose a novel approach using whole model inputs i.e. Average Segment input. Our experiments on PU1K and ABC datasets reveal that patch-based inputs consistently outperform whole model inputs. To understand this, we will delve into factors in feature extraction, and network architecture that influence upsampling results. …
keywords:
@@ Abstract
1 Introduction
@@
The improvement of 3D scanners in 3D sensors, as well as the rapid development of deep learning in computer vision, have led to a gradual shift in the popularity of computer vision technology from 2D image to 3D model. Especially in the direction of autonomous driving and robots that exchange with the physical world, there is a particular focus on three-dimensional vision [1, 2, 3, 4].
@@
Figure 1: The original Mesh models have been transformed into a point cloud through patch based sampling and the Average Segment method proposed in this article.
@@
Vis-a-vis other representations such as Mesh, point clouds are an effective way to represent three-dimensional shapes that are widely used in the aforementioned fields. In addition, the task of reverse engineering CAD models from 3D point clouds is receiving increasing attention [5]. In the process of redesigning the model obtained from reverse engineering, although there are Application Programming Interface (API) available in CAD software [6], if these designs involve reverse engineering and highly customized and personalized designs, CAD APIs are powerless [7]. Therefore, the generation of model point clouds is crucial.
@@
The point cloud upsampling technology based on deep learning has developed with the development of deep learning, extracting multi-resolution features from the initial PU-Net [8] on the basis of pointnet++ [9]. Then, multiple independent multi-layer perceptrons (MLPs) are introduced to generate the required dense point cloud. Moving on to MPU/3PU [10], KNN is used for interpolation and EdgeConv [11] is used to extract features for hierarchical reconstruction through recursive upsampling. Implementing upsampling PU-GAN [12] based on GAN [13] generation architecture, guiding the generator to upsample and approximate the real point cloud. PU-GCN [14] modifies the upsampling module from simple duplicate to Nodeshuffle based on feature extraction in MPU, achieving very good results. And Transformer’s point cloud upsampling network [15, 16]. There are also methods to create coarse point clouds and then refine them by combining global and local methods [17].
@@
The above methods mainly utilize Euclidean spatial features at potential spatial points, upsampling from low resolution to high resolution. The inherent limitations have also been addressed by the work of upsampling point clouds at any scale [18, 19] and capturing local information [20].
@@
Nevertheless, these works still have limitations, which are on the training dataset. Existing point cloud upsampling work [14, 15, 18, 19, 20, 21] without exception used the PU1K dataset [14], which was used based on the methods of MPU and PU-GAN [10, 12]. To obtain training data by randomly sampling the entire 3D model using Patch under conditions of limited computing power, especially limited graphics memory. Undoubtedly, Patch based data input can reduce memory consumption and computational complexity, but it can bring about boundary effects and continuity loss in reverse engineering. It is necessary to explore a new point cloud input method to avoid the above issues. Moreover, the existing point cloud upsampling work focuses on 3D models similar to the PU1K dataset, while neglecting the exploration of point cloud upsampling for parts and CAD. In addition to conducting experiments on the PU1K dataset, this article will also discuss the upsampling effect of point cloud models on workpieces on the ABC dataset [22].
@@
Based on the above issues, in order to explore why numerous point cloud sampling data are still based on Patch input, this paper proposes a new data input method. Unlike dividing point clouds according to Patch, this article takes the point cloud model uniformly and randomly divided and corresponds them one by one as input, which is named as Average Segment (AS). After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN [14] network, and cross validation was performed on both the PU1k and ABC datasets. And during the experiment, the encoder and decoder of PU-GCN [14] were ablated.
@@
To solve the above problems, this paper proposes a new data input method. Unlike dividing point clouds according to patches, this paper uniformly and randomly divides point cloud models and uses them as inputs one by one. After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN [14], and cross validation was performed on both the PU1k [14] and ABC datasets [22]. And during the experiment, the encoder and decoder of PU-GCN [14] were ablated.
@@ 2.1 Overview
2.2 Data Input Based on Patch
@@
2.3 Data Input Based on Average Segment
@@ 3.1 Settings
  • PU1K: This is a point cloud upsampling dataset introduced in PU-GCN [14]. Overall, PU1K consists of 1147 3D models, including 120 3D models from the PU-GAN dataset, as well as 900 different models collected from ShapeNetCore [23]. The test set includes 27 models from PU-GAN and over 100 models from ShapeNetCore. Covering 50 object categories.
  • ABC Dataset: The ABC Dataset [22] is a dataset containing one million computer-aided design (CAD) models, aimed at promoting research on geometric deep learning methods and applications. It is currently one of the largest publicly available datasets for geometric deep learning research.
@@
Figure 2: The PU-GCN network structure used in the experiment is shown in the figure, where (a) is the original PU-GCN structure. (b), (c) and (d) are all architectures used in the ablation experiment to explore the modules that have the greatest impact on upsampling.
3.2 Evaluation Metrics
During the testing process, this article adopted the common practices used in previous work on point cloud upsampling [14, 15, 18, 19, 20, 21]. That is to say, this article will cut the input point cloud into multiple subset Patch and AS of all N points.
In the 4 times upsampling experiment of this article, three types of low resolution to high-resolution sampling were performed. Each test sample has a low resolution point cloud consisting of 256, 512 and 1048 points, and a high-resolution point cloud consisting of 1024,2048 and 8196 points. This article quantitatively evaluates the upsampling performance of PU-GCN under two methods based on three widely used indicators: (i) Chamfer distance (CD) [24], (ii) Hausdorff distance (HD) [25], and (iii) model parameter quantity.
3.3 Results
@@
Table 1: The data input based on the Patch method and AS method is obtained using PU-GCN in PU1K.
@@
Table 2: The data input based on the Patch method and AS method is obtained using PU-GCN in ABC10K.
@@ 3.4 Ablations
Table 3: The results of swapping test sets based on Patch method and AS method on two datasets.
@@
Table 4: Perform ablation experiments on different architectures of PU-GCN using two data input methods.
@@ 3.5 Visualizations
Figure 3: Visualization of PU-GCN with Different Architectures on the PU1K Dataset.
Figure 4: Visualization example on the ABC dataset.
4 Discussion
@@
5 Future Works
PU-GCN model used in this article as an example, the existing Patch based upsampling method is better than AS method we proposed. This may be due to the difference in dataset size between the two methods. In future work, we will delve deeper into the differences in interpretability between Patch and AS methods through representation learning methods. At the same time, due to the similarity between the use of point clouds and the completion of point clouds. We will refer to the method in article  [26] and use encoders for the data input of Patch and AS respectively, in order to merge global and edge features and achieve better results.
6 Conclusion